# Robust $Q$-Learning for finite ambiguity sets

### v2

@@ 5. Numerical Experiments
In this section we provide two numerical examples comparing Algorithm 1 with other robust and non-robust $Q$-learning algorithms. To apply the numerical method from Algorithm 1, we use for all of the following examples a sequence of learning rates defined by $\widetilde{\gamma}_{t}=\frac{1}{1+t}$ for $t\in\mathbb{N}$, a discount factor of $\alpha=0.95$, as well as an $\epsilon$-greedy policy with $\epsilon_{\operatorname{greedy}}=0.1$. Finally, we run [^3] We consider the algorithm as being converged if further iterations do not change the optimal action derived from $Q$-learning, i.e. if $\operatorname*{arg\,max}_{b\in A}Q_{t}(x,b)$ remains constant for further iterations $t$ for all $x\in\mathcal{X}$. This was the case for all presented experiments. Algorithm 1 with $1~{}000~{}000$ iterations on a processor: 13th Gen Intel(R) Core(TM) i7-13700KF 3.40 GHz. Further details of the implementation can be found under https://github.com/CecileDecker/FiniteQLearning.git.
@@ 5.1. Coin Toss
As in [31, Example 4.1] we consider an agent playing a coin toss game.
At each time step $t\in\mathbb{N}$ the agent observes the result of $10$ coins where an outcome of heads corresponds to $1$, and tails corresponds 00. The state $X_{t}$ at time $t\in\mathbb{N}$ is then given by the sum of the $10$ coins value, i.e., we have $\mathcal{X}:=\{0,...,10\}$.
At each time step $t$ the agent can make a bet whether the sum of the next throw strictly exceeds the previous sum (i.e., $X_{t+1}>X_{t}$), or whether it is strictly smaller (i.e., $X_{t+1}<X_{t}$). If the agent is correct, she gets $1\$$, however if the agent is wrong she has to pay $1\$$. The agent also has the possibility not to play. We model this by considering the following reward function:
(5.1) $\mathcal{X}\times A\times\mathcal{X}\ni(x,a,x^{\prime})\mapsto r(x,a,x^{\prime}):=a1\mathrm{l}_{\{x<x^{\prime}\}}-a1\mathrm{l}_{\{x>x^{\prime}\}}-|a|1\mathrm{l}_{\{x=x^{\prime}\}},$
where the possible actions are given by $A:=\{-1,0,1\}$ with $a=1$ corresponding to betting $X_{t+1}>X_{t}$, $a=0$ to not playing, and $a=-1$ to betting $X_{t+1}<X_{t}$. In a next step, we define two different ambiguity sets via [^4] We denote by $\operatorname*{Bin}(n,p)$ a binomial distribution with $n$ number of trials and $p$ being the probability of success in a single trial.
@@
and we aim at comparing Algorithm 1 with the Algorithm from [31] which however does not allow to build the same assymmetric sets as in (5.2) and (5.3). To build comparative ambiguity sets with Wasserstein uncertainty we define ambiguity sets such that all probability measures from $\mathcal{P}^{1}$ and $\mathcal{P}^{2}$ are contained, respectively. This leads to [^5] Here $W_{1}$ denotes the Wasserstein-$1$ distance, compare for more details, e.g., [47].
(5.4) $\mathcal{P}^{3}(x,a):=$$\left\{\mathbb{P}\in\mathcal{M}_{1}(\mathcal{X})~{}\middle|~{}W_{1}(\mathbb{P},\operatorname*{Bin}(10,0.5))\leq 1\right\}$
(5.5) $\mathcal{P}^{4}(x,a):=$$\left\{\mathbb{P}\in\mathcal{M}_{1}(\mathcal{X})~{}\middle|~{}W_{1}(\mathbb{P},\operatorname*{Bin}(10,0.5))\leq 2\right\}$
which are Wasserstein-balls around the reference probability $\operatorname*{Bin}(10,0.5)$ with radii $1$ and $2$, respectively. As one can see, e.g., from [29, Equation (4.2)], $\mathcal{P}^{3}$ is the smallest Wasserstein-ball around the reference measure that contains $\operatorname*{Bin}(10,0.6)$, whereas $\mathcal{P}^{4}$ is the smallest Wasserstein-ball containing $\operatorname*{Bin}(10,0.3)$.
We depict the trained actions and the associated convergence time (duration of running $1~{}000~{}000$ iterations of the algorithm) in Table 1.
$X_{t}$ 0 1 2 3 4 5 6 7 8 9 10 Convergence time
[…]
Table 1. The trained actions $a_{t}^{\mathcal{P}^{1}}$, $a_{t}^{\mathcal{P}^{2}}$, $a_{t}^{\mathcal{P}^{3}}$, $a_{t}^{\mathcal{P}^{4}}$, $a_{t}^{\rm non-robust}$, in dependence of the realized state $X_{t}$
We next test the profit of the actions described in Table 1 by playing $100~{}000$ rounds of the game according to the trained policies $a_{t}^{\mathcal{P}^{1}}$, $a_{t}^{\mathcal{P}^{2}}$ ,$a_{t}^{\mathcal{P}^{3}}$ and $a_{t}^{\mathcal{P}^{4}}$. For simulating the $100~{}000$ rounds we assume an underlying binomial distribution $\mathbb{P}=\operatorname*{Bin}(10,p_{\rm true})$ with a fixed probability $p_{\rm true}$ for heads which we vary from $0.1$ to $0.9$. We depict the cumulative profits of the considered actions in Table 3.
@@
Robust, $\mathcal{P}^{1}$ $-31240$ $-19121$ $-4283$ $15161$ 3048430484{30484}30484 $\mathbf{29857}$ 1321213212{13212}13212 $-10339$ ${-29843}$
Robust, $\mathcal{P}^{2}$ $\mathbf{-24424}$ $\mathbf{5231}$ $\mathbf{21051}$ $\mathbf{24613}$ 2299822998{22998}22998 $11715$ $-4984$ $-18896$ $-31185$
Robust, $\mathcal{P}^{3}$ $-30003$ $-10859$ $9843$ $21900$ $25395$ $22587$ $9653$ $-10656$ $-30301$
Robust, $\mathcal{P}^{4}$ $-24528$ $4263$ $16771$ $13415$ $9925$ $13366$ $\mathbf{17345}$ $\mathbf{4731}$ $\mathbf{-24452}$
Non-Robust $-31101$ $-18251$ $-521$ $23175$ $\mathbf{35244}$ $23227$ $-1387$ $-18421$ $-31024$
@@
We first observe that only in the case $p_{\rm true}=0.5$ the non-robust strategy is the best performing strategy. The Wasserstein-strategies are outperformed by the robust strategies taking into account either $\mathcal{P}_{1}$ or $\mathcal{P}_{2}$ in all cases for which $p_{\rm true}\leq 0.6$, in particular for the cases $p_{\rm true}=0.3,0.6$ which are by construction contained in the respective ambiguity sets.
@@
If an applicant however simply has a best guess of a reference measure and considers the cases $p_{\rm true}=0.4$ and $p_{\rm true}=0.6$ both as equally possible, then it is advisable to pursue the Wasserstein-approach outlined in [29].
5.2. Stock investing
This example is a variant of [31, Example 4.3] showcasing the use of tailored ambiguity sets which are not possible to construct and use in the same manner with Wasserstein or Entropic ambiguity sets as they are considered in [7], [23], and [48].
We consider the problem of optimally investing in the stock market, and, to this end, encode the return of a stock by $2$ numeric values: either the return is positive ($1$) or negative ($-1$). The space of numerically encoded returns is therefore given by: $T:=\{-1,1\}$. At each time step $t$ the agent can choose to buy the stock or not, or to short sell the stock. The action $a_{t}$ represent this investment decision, buying the stock is encoded by $1$ and not buying it by 00, whereas short-selling corresponds to $-1$, i.e., we have $A:=\{-1,0,1\}$.
The agent’s investment decision should depend not only on the most recent return but naturally will depend on the current market situation. Therefore the agent relies her investment decision on the last $h=5$ returns. Hence, we consider the state space
$\mathcal{X}:=T^{h}=\{-1,1\}^{h}.$
@@ 5.2. Stock investing
(5.6) $\mathcal{X}\times A\times\mathcal{X}\ni(x,a,x^{\prime})\mapsto r(x,a,x^{\prime}):=a\cdot x^{\prime},$
i.e., we reward a correct investment decision and penalize an incorrect investment decision.
Now, to train the agent, we construct an estimated reference measure in the following way. For the state transition, we consider the historic evolution of the (numerically encoded) returns of the underlying stock. This time series is denoted by $(R_{j})_{j=1,...,N}\subset T^{N}$ for some $N\in\mathbb{N}$. We then define for some small $\gamma>0$ a map [^7] We only employ $\gamma$ to avoid division by 00, hence $\gamma$ can be chosen arbitrarily small. $\mathcal{X}\times A\ni(x,a)\mapsto\widetilde{\mathbb{P}}(x,a):=\sum_{i\in T}p_{i}(x)\cdot\delta_{\{i\}}\in\mathcal{M}_{1}(T)$ where for $x\in\mathcal{X}$, $i\in T$ we define:
(5.7) $p_{i}(x):=\frac{{\widetilde{p}_{i}}(x)+\frac{\gamma}{4}}{\gamma+\sum_{j\in T}\widetilde{p}_{j}(x)}\in[0,1],$
with
(5.8) $\widetilde{p}_{i}(x):=\sum_{j=1}^{X-h+1}1\mathrm{l}_{\{(\pi(x),i)=(R_{j},...,R_{j+h-1})\}},$
and where $\mathbb{R}^{h}\ni(x_{1},\dots,x_{h})\mapsto\pi(x_{1},\dots,x_{h}):=(x_{2},\dots,x_{h})\in\mathbb{R}^{h-1}$ denotes the projection onto the last $h-1$ components. This means the data-driven construction of the measure ~$\widetilde{\mathbb{P}}(x,a)$ relies on the relative frequency of the sequence $(\pi(x),i)$ in the time series of past realized returns $(R_{j})_{j=1,...,N}$ for $i\in\{-1,1,\}$: the more often the market situation $(\pi(x),i)\in\mathcal{X}$ was observed in the training data $(R_{j})_{j=1,\dots,N}$ the higher the value $\widetilde{p}_{i}$. Equation (5.7) is then applied to convert the frequencies $\widetilde{p}_{i}$ to probabilities $p_{i}$. Then, the state transition of the reference measure is partially deterministic ($h-1$ components of subsequent states coincide) and is of the form
(5.9) $X\times A\ni(x,a)\mapsto\mathbb{P}(x,a)=\delta_{\pi(x)}\otimes\widetilde{\mathbb{P}}(x,a)\in\mathcal{M}_{1}(\mathcal{X}).$
To construct the probability measure from (5.9) using real data, we consider daily returns of the stock of Alphabet Inc (previously known as Google) in the time period from beginning of $2010$ until $2024$, where we depict the cumulative sum of the signs of the returns in Figure 1.
@@
We split the data into a training period ranging from January $2010$ until December $2021$ and two testing period starting directly thereafter until February $2023$ and January $2024$, respectively. We then first construct a measure $\mathbb{P}_{1}$ according to (5.9) by using all training data. To construct an ambiguity set $\mathcal{P}$ we isolate specific periods of the training data where the market behaved very volatile (Training Period 1), bearish (Training Period 2) or bullish (Training Period 3). Including the accompanying probability measures $\mathbb{P}_{2},\mathbb{P}_{3},\mathbb{P}_{4}$ in an ambiguity set $\mathcal{P}:=\{\mathbb{P}_{1},\mathbb{P}_{2},\mathbb{P}_{3},\mathbb{P}_{4}\}$ then accounts for the fear of the agent to not trade profitably in these difficult market scenarios. The choice of these specific training periods is depicted in Figure 1. Using Algorithm 1, we then train different agents using the ambiguity set $\mathcal{P}$ and using the non-robust state transitions $\mathbb{P}_{i}$, $i=1,2,3,4$, respectively. Moreover, we consider a trend following strategy (i.e., always invest according to the sign of the last return) as well as a simple buy-and-hold strategy. The results (mean average reward per trade) of the different strategies evaluated in the test period are shown in Table 3, and they show that using a robust strategy indeed enables to avoid losses in both test periods. Note moreover that in the bearish test period 1 the robust strategy in particular outperforms the non-robust agent trained on all data which is certainly due to the inclusion of $\mathbb{P}_{2}$ and $\mathbb{P}_{3}$ in the ambiguity set $\mathcal{P}$ which are the best performing strategies in this period. Also note that in test period 2 it becomes visible that the robust approach due to its construction of a worst-case approach is better suited to avoid losses than to profit from bearish market scenarios.
Test Period 1 Test Period 2
Robust ($\mathcal{P}$) 0.0205 0.0521
Non-Robust: All Data ($\mathbb{P}_{1}$) -0.0274 0.0729
Non-Robust: Training Period 1 ($\mathbb{P}_{2}$) 0.0377 0.0885
Non-Robust: Training Period 2 ($\mathbb{P}_{3}$) 0.0685 0.1042
Non-Robust: Training Period 3 ($\mathbb{P}_{4}$) -0.0034 0.1458
Trend following -0.0137 0.0781
Buy and hold -0.0822 0.1146
Table 3. A comparison of the average reward per trade in the two test periods depicted in Figure 1 among different strategies. The best-performing strategy is highlighted with bold letters.

### v3

@@ 6. Numerical Experiments
In this section we provide two numerical examples comparing Algorithm 1 with other robust and non-robust $Q$-learning algorithms. To apply the numerical method from Algorithm 1, we use for all of the following examples a sequence of learning rates defined by $\widetilde{\gamma}_{t}=\frac{1}{(1+t)^{0.6}}$ for $t\in\mathbb{N}$, a discount factor of $\alpha=0.95$, as well as an $\epsilon$-greedy policy with $\varepsilon$ decreasing over number of iterations, starting with $1$, and decreasing linearly to $0.05$. All the reported experiments are being executed with 10000001~000~000 iterations on a 11th Gen Intel(R) Core(TM) i7-1165G7 @ 2.80GHz (2.80 GHz) - processor. Further details of the implementation can be found under https://github.com/juliansester/FiniteQLearning.
@@ 6.1. Coin toss
Inspired by [38, Example 4.1], we consider an agent playing a coin toss game.
At each time step $t\in\mathbb{N}$ the agent observes the result of $10$ coins where an outcome of heads corresponds to $1$, and tails corresponds to 0. The state $X_{t}$ at time $t\in\mathbb{N}$ is then given by the sum of the $10$ coins value, i.e., we have $\mathcal{X}:=\{0,\cdots,10\}$.
At each time step $t$ the agent can make a bet whether the sum of the next throw strictly exceeds the previous sum (i.e., $X_{t+1}>X_{t}$), or whether it is strictly smaller (i.e., $X_{t+1}<X_{t}$). If the agent is correct, she gets $1\mathdollar$, in contrast if the agent’s bet is wrong she has to pay $1\mathdollar$. The agent also has the possibility not to play. We model this by considering the following reward function:
(6.1) $\mathcal{X}\times A\times\mathcal{X}\ni(x,a,x^{\prime})\mapsto r(x,a,x^{\prime}):=a1\hskip-3.0pt\mathrm{l}_{\{x<x^{\prime}\}}-a1\hskip-3.0pt\mathrm{l}_{\{x>x^{\prime}\}}-|a|1\hskip-3.0pt\mathrm{l}_{\{x=x^{\prime}\}},$
where the possible actions are given by $A:=\{-1,0,1\}$ with $a=1$ corresponding to betting $X_{t+1}>X_{t}$, $a=0$ to not playing, and a=1a=-1 to betting $X_{t+1}<X_{t}$. In a next step, we define two different ambiguity sets via [^4] We denote by $\operatorname*{Bin}(n,p)$ a binomial distribution with $n$ number of trials and $p$ being the probability of success in a single trial.
@@
and we aim at comparing Algorithm 1 with the Algorithm from [38] which does not directly allow to build the same asymmetric sets as in (6.2) and (6.3). To build comparative ambiguity sets with Wasserstein uncertainty we define ambiguity sets such that all probability measures from $\mathcal{P}^{1}$ and $\mathcal{P}^{2}$ are contained, respectively. This leads to [^5] Here $W_{1}$ denotes the Wasserstein-$1$ distance, compare for more details, e.g., [58].
(6.4) $\mathcal{P}^{3}(x,a):=$ $\left\{\mathbb{P}\in\mathcal{M}_{1}(\mathcal{X})~\middle|~W_{1}(\mathbb{P},\operatorname*{Bin}(10,0.5))\leq 1\right\},$
(6.5) $\mathcal{P}^{4}(x,a):=$ $\left\{\mathbb{P}\in\mathcal{M}_{1}(\mathcal{X})~\middle|~W_{1}(\mathbb{P},\operatorname*{Bin}(10,0.5))\leq 2\right\},$
which are Wasserstein-balls centered at the reference probability $\operatorname*{Bin}(10,0.5)$ with radii $1$ and $2$, respectively. As one can see, e.g., from [39, Equation (4.2)], $\mathcal{P}^{3}$ is the smallest Wasserstein-ball around the reference measure that contains $\operatorname*{Bin}(10,0.6)$, whereas $\mathcal{P}^{4}$ is the smallest Wasserstein-ball containing $\operatorname*{Bin}(10,0.3)$, but, by construction they contain many additional distributions (e.g. $\operatorname*{Bin}(10,p)$ with intermediate $p$). This highlights the unavoidable over-coverage induced by metric balls.
We train optimal actions for $\mathcal{P}^{1},\mathcal{P}^{2}$ by Algorithm 1, and optimal actions for $\mathcal{P}^{3},\mathcal{P}^{4}$ using the Algorithm from [39]. Upon training, we depict the trained actions in Table 1.
state $X_{t}$ 0 1 2 3 4 5 6 7 8 9 10
$a_{t}^{\mathcal{P}^{1}}(X_{t})$ 11 11 11 11 11 0 0 1-1 1-1 1-1 1-1
$a_{t}^{\mathcal{P}^{2}}(X_{t})$ 11 11 11 0 0 0 1-1 1-1 1-1 1-1 1-1
$a_{t}^{\mathcal{P}^{3}}(X_{t})$ 11 11 11 11 0 0 0 1-1 1-1 1-1 1-1
$a_{t}^{\mathcal{P}^{4}}(X_{t})$ 11 11 11 0 0 0 0 0 1-1 1-1 1-1
$a_{t}^{\rm non-robust}(X_{t})$ 11 11 11 11 11 0 1-1 1-1 1-1 1-1 1-1
Table 1. The trained actions $a_{t}^{\mathcal{P}^{1}}$, $a_{t}^{\mathcal{P}^{2}}$, $a_{t}^{\mathcal{P}^{3}}$, $a_{t}^{\mathcal{P}^{4}}$, and $a_{t}^{\rm non-robust}$ in dependence of the realized state $X_{t}$
We test the profitability of the actions described in Table 1 by playing 100000100~000 rounds of the game according to the trained policies $a_{t}^{\mathcal{P}^{1}}$, $a_{t}^{\mathcal{P}^{2}}$ $a_{t}^{\mathcal{P}^{3}}$ and $a_{t}^{\mathcal{P}^{4}}$. For simulating the 100000100~000 rounds we assume an underlying binomial distribution $\mathbb{P}=\operatorname*{Bin}(10,p_{\rm true})$ with a fixed probability $p_{\rm true}$ for heads which we vary from $0.1$ to $0.9$. We depict the profits of the considered actions in Table 3 and in Figure 1.
Moreover, in Figure 2 we illustrate the convergence of $\max_{(x,a)\in\mathcal{X}\times A}|Q_{t}(x,a)-Q^{*}(x,a)|$ to 0 as the number of itrations $t$ increases. To this end, we study the two ambiguity sets $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ in the two cases where the transition probabilities are entirely known and are being learned from samples, respectively. We see that in both cases the algorithm converges after a reasonable number of iterations ($\approx 100,000$) up to a small tolerance, while the convergence without known probabilities is slower, accounting for the fact that there is an additional approximation error due to the mismatch between learned probability and true probability, compare also Theorem 3.4.
With the mentioned computational setup, the compute time for $500,000$ iterations was around $3$ minutes for all of the discussed variants.
@@
Robust, $\mathcal{P}^{1}$ 31240-31240 19121-19121 4283-4283 $15161$ 30484{30484} $\mathbf{29857}$ 13212{13212} 10339-10339 29843{-29843}
Robust, $\mathcal{P}^{2}$ $\mathbf{-24424}$ $\mathbf{5231}$ $\mathbf{21051}$ $\mathbf{24613}$ 22998{22998} $11715$ 4984-4984 18896-18896 31185-31185
Robust, $\mathcal{P}^{3}$ 30003-30003 10859-10859 $9843$ $21900$ $25395$ $22587$ $9653$ 10656-10656 30301-30301
Robust, $\mathcal{P}^{4}$ 24528-24528 $4263$ $16771$ $13415$ $9925$ $13366$ $\mathbf{17345}$ $\mathbf{4731}$ $\mathbf{-24452}$
Non-Robust 31101-31101 18251-18251 521-521 $23175$ $\mathbf{35244}$ $23227$ 1387-1387 18421-18421 31024-31024
@@
Figure 1. Average per step reward of the game described in Section 6.1 in dependence of different trained strategies (described in Table 1) and of the probability distribution $P=\operatorname*{Bin}(10,p_{\rm true})$ of the simulated underlying process, where $p_{\rm true}$ is indicated on the $x$-axis.
Figure 2. The figure shows the convergence of $\max_{(x,a)\in\mathcal{X}\times A}|Q_{t}(x,a)-Q^{*}(x,a)|$ for the two ambiguity sets $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ for the cases when the transition probabilities are explicitly learned and are being learned from samples, respectively.
For the base case $p_{\rm true}=0.5$ the non-robust strategy is the best performing strategy. The Wasserstein-strategies are outperformed by the robust strategies taking into account either $\mathcal{P}_{1}$ or $\mathcal{P}_{2}$ in all cases for which $p_{\rm true}\leq 0.6$, in particular for the cases $p_{\rm true}=0.3,0.6$ which are by construction contained in the respective ambiguity sets.
@@
If an applicant however simply has a best guess of a reference measure and considers the cases $p_{\rm true}=0.4$ and $p_{\rm true}=0.6$ both as equally possible, then it is advisable to pursue the Wasserstein-approach outlined in [39]. To summarize the non-robust strategy dominates exactly at the reference model, the finite-robust strategy dominate exactly on the respective ambiguity set, and the Wasserstein robust strategy dominate only when ambiguity is genuinely diffuse.
Example 6.1 thus illustrates how the choice of the ambiguity set fundamentally affects the structure of the optimal policy and the resulting performance. When the ambiguity set is specified as a finite collection of candidate transition kernels, as in $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$, the robust $Q$-learning algorithm yields policies that are tailored to protect against a small number of explicitly modelled adverse scenarios. In particular, the resulting strategies exhibit asymmetric betting regions and early abstention in states where the worst-case model renders the expected payoff close to zero.
In contrast, Wasserstein ambiguity sets such as $\mathcal{P}_{3}$ and $\mathcal{P}_{4}$, which are defined as metric balls around a reference distribution, necessarily include a continuum of intermediate distributions that were not explicitly intended to be protected against. This leads to more conservative policies, characterized by wider neutral regions in which the agent refrains from betting. While this conservatism provides robustness against diffuse model misspecification, it may be suboptimal when the modeller has prior knowledge that uncertainty is concentrated on a small number of specific alternatives.
The numerical results confirm this distinction. When the true data-generating process coincides with one of the distributions contained in the finite ambiguity sets (e.g. $p_{\mathrm{true}}\in\{0.5,0.6\}$ for $\mathcal{P}_{1}$), the corresponding finite-robust strategy consistently outperforms both the Wasserstein-robust and the non-robust strategies. Conversely, when uncertainty is genuinely diffuse and symmetric around the reference measure, Wasserstein-based robustness provides a reasonable compromise.
Overall, this example highlights that finite ambiguity sets enable a form of scenario-based robustness, whereas Wasserstein balls implement neighbourhood-based robustness. The former is preferable when the modeller can identify a small set of plausible adverse scenarios, while the latter is appropriate when uncertainty is best described by deviations around a reference model without clear directional structure.
6.2. Portfolio allocation
This example is a variant of [38, Example 4.3] showcasing the use of tailored asymmetrical ambiguity sets which are not possible to construct and use in the same manner with Wasserstein or Entropic ambiguity sets as they are considered in [8], [29], and [59].
We consider the problem of optimally investing in the stock market, and, to this end, encode the return of a stock by $2$ numeric values: either the return is positive ($1$) or negative (1-1). The space of numerically encoded returns is therefore given by $T:=\{-1,1\}$. At each time step $t$ the decision maker can choose to buy the stock or not, or to short sell the stock. The action $a_{t}$ represent this investment decision, where buying the stock is encoded by $1$ and not buying it by 0, whereas short-selling corresponds to 1-1, i.e., we have $A:=\{-1,0,1\}$.
The agent’s investment decision should depend not only on the most recent return but naturally will depend on the current market situation reflecting autocorrelation of financial time series. Therefore the agent relies her investment decision on the last $h=5$ returns and on the previous position. Hence, we consider the state space
$\mathcal{X}:=T^{h}\times A=\{-1,1\}^{h}\times\{-1,0,1\}.$
@@ 6.2. Portfolio allocation
$r:\mathcal{X}\times A\times\mathcal{X}$ $\rightarrow\mathbb{R}$
$\bigg((x_{-h+1},\cdots,x_{0},a_{0}),a_{1},(x_{-h+2},\cdots,x_{1},a_{1})\bigg)$ $\mapsto a_{1}\cdot x_{1}-\eta\cdot|a_{1}-a_{0}|,$
i.e., we reward a correct investment decision and penalize an incorrect investment decision. Additionally, the term $\eta\cdot|a_{1}-a_{0}|$ penalizes position changes, accounting for transaction costs. In our experiments we set $\eta=0.1$.
Now, to train the agent, we construct market simulators in the following way. We consider the historical evolution of the (numerically encoded by 1-1 and $1$) daily returns of the underlying stock. This time series is denoted by $(R_{j})_{j=1,...,N}\subset T^{N}$ for some $N\in\mathbb{N}$ labeling the length of the time series. Given some $x=(x_{-h+1},\cdots,x_{0},a_{0})\in\mathcal{X}$ we then sample uniformly the next return from all historical returns sequences with coinciding history [^7] If no such sequence exist we sample uniformly from $T$. $(x_{-h+1},\cdots,x_{0})$. Then, the state transition of the reference measure is partially deterministic (h1h-1 components of subsequent states coincide) and the next state is of the form
$(x_{-h+2},\cdots,x_{0},y,a_{1})\in\mathcal{X}$
where $y$ is the sampled return and $a_{1}$ the executed action.
To apply this construction to real data, we consider daily returns of the stocks of Alphabet, Nvidia, Apple, Microsoft, Amazon, Broadcom, JP Morgan (Tickers: ”AAPL”, ”MSFT”, ”AMZN”, ”GOOGL”,”AVGO”,”JPM”) in the time period from January $2010$ until December $2025$, where we depict the cumulative sum of the signs of the returns in Figure 3.
@@
We split the data into a long training period ranging from January $2010$ until December $2021$ and three testing period starting directly thereafter until February $2023$, January $2024$, and November $2025$, respectively. To construct an ambiguity set $\mathcal{P}$ we isolate specific periods of the training data where the market behaved very volatile (Training Period 1), bearish (Training Period 2) or bullish (Training Period 3). These periods vary from stock to stock and are determined by checking the whole training period for that sub-period of $250$ days with 1) the largest standard deviation, 2) the largest increase, and 3) the largest decrease. We denote the probability measures corresponding to these sub-periods by $\mathbb{P}_{1},\mathbb{P}_{2},\mathbb{P}_{3}$, respectively. Moreover, we construct a market simulator (with distribution denoted by $\mathbb{P}_{4}$) by using all training data. Including the accompanying probability measures $\mathbb{P}_{1},\mathbb{P}_{2},\mathbb{P}_{3},\mathbb{P}_{4}$ in an ambiguity set $\mathcal{P}:=\{\mathbb{P}_{1},\mathbb{P}_{2},\mathbb{P}_{3},\mathbb{P}_{4}\}$ then accounts for the fear of the agent to not trade profitably in all considered (difficult) market scenarios, and in normal periods ($\mathbb{P}_{4}$). The choice of these specific training periods is depicted in Figure 3. Using Algorithm 1, we then train different agents using the the market simulators corresponding to the ambiguity set $\mathcal{P}$ and the market simulators for the non-robust state transitions $\mathbb{P}_{i}$, $i=1,2,3,4$, respectively. We emphasize that to use Algorithm 1 we do not need to know $\mathbb{P}_{1},\mathbb{P}_{2},\mathbb{P}_{3},\mathbb{P}_{4}$ explicitly but only need to be able to sample from these measures.
As a benchmark for evaluation, we consider a simple buy-and-hold strategy. The results (mean average reward per trade across all tickers) of the different strategies evaluated in the test period are shown in Table 3, and they show that using a robust strategy indeed enables to avoid losses in all test periods. Note moreover that in the bearish test period 1 the robust strategy in particular outperforms the non-robust agent trained on all data which is likely due to the inclusion of $\mathbb{P}_{3}$ in the ambiguity set $\mathcal{P}$ which is the best performing strategies in this period. Also note that in test periods 2 and 3 it becomes visible that the robust approach due to its construction of a worst-case approach is better suited to avoid losses than to profit from bearish market scenarios.
Test Period 1 Test Period 2 Test Period 3
Days 299 199 500
Negative Returns 1078 635 1573
Positive Returns 1015 758 1927
Average Rewards
Robust ($\mathcal{P}$) 0.0005 0.0168 0.0069
Non-Robust: Training Period 1 ($\mathbb{P}_{1}$) -0.0460 0.0645 0.0150
Non-Robust: Training Period 2 ($\mathbb{P}_{2}$) -0.0436 0.0733 0.0901
Non-Robust: Training Period 3 ($\mathbb{P}_{3}$) 0.0368 -0.0462 -0.0827
Non-Robust: All Data ($\mathbb{P}_{4}$) -0.0397 0.0469 0.0559
Buy and hold -0.0392 0.0989 0.1025
Table 3. A comparison of the average reward per trade in the three test periods depicted in Figure 3 among different strategies and across all tickers. The best-performing strategy is highlighted with bold letters.